OpenStudy (gabylovesyou):

Calculate the area of triangle CDE with altitude EF, given C (3, −2), D (−1, 2), E (2, 3), and F (0, 1). 4 square units 6.2 square units 8 square units 8.7 square units

jagr2713 (jagr2713):

i think it is C

jagr2713 (jagr2713):

A=b*h/2

OpenStudy (ranga):

CD is the base of the triangle. Find the length of CD using the distance formula. EF is the altitude of the triangle. Find the length of EF using the distance formula. Area = 1/2 * CD * EF

jagr2713 (jagr2713):

so its c right

OpenStudy (anonymous):

do you mind saying the answer because her OS is not working and she wants me to tell her

jagr2713 (jagr2713):

Its C

OpenStudy (anonymous):

thanks guys

jagr2713 (jagr2713):

You're welcome

OpenStudy (e.mccormick):

Check your numbers.

OpenStudy (e.mccormick):

\(\dfrac{1}{2}\cdot \sqrt{((-1)-3)^2+(2-(-2))^2}\cdot \sqrt{(2-0)^2+(3-1)^2}\) Always best to write it out. Then you can be sure yourself, rather than just relying on others.

OpenStudy (anonymous):

i know its not my question but gaby cant get on and i dont know how to do it. i had to text her the answer